Make New Entry,
Make Followup Entry
User name J. Wexler
Log entry time 08:08:43 on August28,2009
Entry number 286519
keyword=Shift summary - Owl 8/28
Shift summary - Owl 8/28:
12AM - Request 5uA on the optics target.
12:25- Attempted to move to cryo target, but control issue leads us to go to the BeO target to see if we still have alignment after cooling target.
12:38- Target control fixed by restarting GUI. Visually centered beam on target at -1.8, 2.2 tune on BPM B.
1:00 - Checked raster, found 2mm x 2mm raster. Also going to hydrogen target (runs 25464-5, 4382-4).
1:45 - Started HAPPEX detector checks. Put sieve in. Start at 5uA (eventually increased to 10 uA). (Runs 4385, 24566)
2:15 - Switch to optics target, raster off. (Runs 24567-8, 4386-8)
2:35 - Ramp current to 70 uA.
3:10 - Take out sieve, run on Al dummy (runs 24570, 4389).
3:40 - Noted differing rates on the two arms, so we equalized the voltages on the S0 triggers on both arms at -1650 V.
3:50 - Thick Ta target in. (runs 24570-5, 4390-1)
~4:10 - Lowered left arm momentum to 1.5 GeV/c (exact time unknown)
4:55 - Changed to hydrogen target. Runs 4395, 24579.
5:30 - Stopped HAPPEX detector checks, changed right arm momentum to 1.5 GeV/c.
5:40 - Redid carbon hole tests w/ spot++ because of the effects of the target being cold.
6:30 - Found correct position tune of BPM B x,y = -2.2, 2.2. Begin harp scans.
7:10 - Finished tuning harp scans. Intrinsic size on targets approximately 180 x 180. Begin raster mapping.
7:30 - Raster mapped. Decided to use 2.5 x 3 mm raster (MCC 5x5) on hydrogen target.
7:40 - Changed to hydrogen target, but realized right dipole still ramping, preventing right arm checks. Proceed to change to empty target from hydrogen target.